# Fluctuations for non-Hermitian dynamics

### v1

@@ 4 Independence of singular values
${\rm d}f_{s}=m_{s}(z)\partial_{z}f_{s}{\rm d}s+\frac{1}{2N}\sum_{|k|\leqslant N}\frac{v_{k}(s)({\rm d}\langle B_{k}\rangle_{s}-{\rm d}s)}{(x_{k}(s)-z)^{3}}-\frac{1}{\sqrt{2N}}\sum_{|k|\leqslant N}\frac{v_{k}(s)}{(z-x_{k}(s))^{2}}{\rm d}B_{k}(s).$ (4.66)
@@
$\begin{split}{\rm d}g_{u\wedge\tau}(z)=(m_{u}(z_{t-u})-\widetilde{m}_{u}(z_{t-u}))\partial_{z}f_{u}(z_{t-u}){\rm d}({u\wedge\tau})+\frac{1}{2N}\sum_{i}\frac{({\rm d}\langle B_{i}\rangle_{u}/{\rm d}u-1)v_{i}(u)}{(x_{i}(u)-z_{t-u})^{3}}{\rm d}(u\wedge\tau)\\ -\frac{\operatorname{O}(1)}{\sqrt{N}}\sum_{k}\frac{{v}_{k}({u})}{(z_{t-u}-x_{k}(u))^{2}}{\rm d}B_{k}({u\wedge\tau}).\end{split}$ (4.68)
@@
where the above integral is due to points $x_{i}(u)\geqslant G/10$, and use the local law in the domain $G/10<x_{i}<G/2$, the upper bound $x_{i}<C$, and the maximum principle bound $|v_{i}|\leqslant 1$. The above integral contributes to at most
@@
$\frac{1}{N}\sum_{i}\frac{({\rm d}\langle B_{i}\rangle_{u}/{\rm d}u-1)v_{i}(u)}{(x_{i}(u)-z_{t-u})^{3}}=\partial_{zz}\frac{1}{2N}\sum_{i}\frac{({\rm d}\langle B_{i}\rangle_{u}/{\rm d}u-1)v_{i}(u)}{x_{i}(u)-z_{t-u}}=\frac{1}{N}\sum_{i\geqslant 1}\partial_{zz}\frac{\operatorname{O}(1)x_{i}v_{i}(u)}{x_{i}(u)^{2}-z_{t-u}^{2}}.$ (4.73)
@@
$M_{s}$ $:=\operatorname{Im}\int_{0}^{s}\frac{1}{\sqrt{N}}\sum_{k}\frac{{v}_{k}({u})}{(z_{t-u}-x_{k}(u))^{2}}{\rm d}B_{k}(u\wedge\tau)$
@@
$\sup_{0\leqslant s\leqslant t}|M_{s}|^{2}\leqslant\varphi^{1/10}\int_{0}^{t}\frac{\eta_{t-u}^{2}}{N}\left(\sum_{k}\left|\frac{E-x_{k}(u)}{|z_{t-u}-x_{k}(u)|^{4}}+\frac{E+x_{k}(u)}{|z_{t-u}+x_{k}(u)|^{4}}\right|{v}_{k}({u})\right)^{2}{\rm d}(u\wedge\tau).$ (4.74)
With (4.72) and the local law ($\tau\leqslant\tau_{0}$) the above sum is bounded with (we omit the terms corresponding to $x_{i}>G/10$ as they are shown to be negligible thanks to the local law and the simple bound $|v_{i}|\leqslant 1$)
[…]
@@
For any $t\in[\varphi^{100}/N,1]$ and $|i|<N$, uniformly in $\alpha$ we have, with overwhelming probability,
@@
We have $|v_{i}(0)|\leqslant\varphi(v_{i}^{(1)}(0)+v_{i}^{(2)}(0))$ from (4.58); from Lemma 4.18 (i), this implies that for any $t\geqslant 0,i\geqslant 1$ we have $|v_{i}(t)|\leqslant\varphi(v_{i}^{(1)}(t)+v_{i}^{(2)}(t))$. From Equation (4.71) and Lemmas 4.20 and 4.21, we have
@@ 5 Resolvent fluctuations from the dynamics noise
$-\frac{1}{m^{z}}=\mathrm{i}\eta+m^{z}-\frac{|z|^{2}}{\mathrm{i}\eta+m^{z}},\qquad\eta{\rm Im}[m^{z}]>0.$
@@
In this section we analyze the evolution of the resolvent $G_{t}^{z}(\mathrm{i}\eta):=(W_{t}-Z-\mathrm{i}\eta)^{-1}$ along the flow (1.5). Here we perform the analysis in the regime $|z|^{2}\leqslant 1-c$, for some small fixed $c>0$, since this is the regime of interest of Proposition 3.3. The main result of this section is a decomposition theorem for $I_{\eta_{c}}^{T}(f,t)$ into the sum of two terms: one that depends only on the initial condition and one that is a martingale (see Proposition 5.1). Before stating this result we introduce the notation
@@
In order to control the evolution of the resolvent along the flow (5.3) we consider the characteristics [^7] Our convention here is that the characteristics move upwards as in (4.37), contrary to (4.63). (cf. [CipErdSch2022bis, Eq. (5.4)], [CipErdXu2023bis, Eq. (5.4)]):
$\partial_{t}\eta_{t}=-{\rm Im}m^{z_{t}}(\mathrm{i}\eta_{t})-\frac{\eta_{t}}{2},\qquad\quad\partial_{t}z_{t}=-\frac{z_{t}}{2}.$ (5.4)
Note that along the characteristics we have
@@
with very high probability, where we used that $|\langle G_{t}^{2}-M_{t}^{\prime}\rangle|\lesssim N^{\xi}(N\eta_{t}^{2})^{-1}$ for any arbitrary small $\xi>0$ by [CipErdSch2022bis, Theorem 3.3] for $z_{1}=z_{2}$ and $\eta_{1}=\eta_{2}$.
@@
Before presenting these proof we comment on their relation with the proof of [CipErdSch2023, Proposition 3.3]. In the reminder of this section the local law [CipErdSch2022bis, Theorem 5.2] is a fundamental input to compute the deterministic approximation of product of resolvents of the form $G_{s}^{z_{1}}G_{s}^{z_{2}}$, i.e. this local law is used for any fixed time $0\leqslant s\leqslant t$, but only for resolvents evaluated at the same time. On the other hand, the CLT for resolvents [CipErdSch2023, Proposition 3.3] is used only at time $s=0$ (see (5.9) below) to compute the correlation of the initial condition.
@@
From now on $c>0$ is a small fixed constant that may change from line to line. By [CipErdSch2023, Proposition 3.3], for $\eta_{i}\geqslant n^{-1+\delta_{1}}$, it follows that
$\begin{split}&\mathbf{E}\mathrm{Tr}\big{[}G_{0}^{v(z_{1})}(\mathrm{i}\eta_{1})-\mathbb{E}G_{0}^{v(z_{1})}(\mathrm{i}\eta_{1})\big{]}\mathrm{Tr}\big{[}G_{0}^{z_{2}}(\mathrm{i}\eta_{2})-\mathbb{E}G_{0}^{z_{2}}(\mathrm{i}\eta_{2})\big{]}\\ &=\partial_{\eta_{1}}\partial_{\eta_{2}}\log\Big{[}1+e^{-t}(|v(z_{1})z_{2}|^{2}u_{1}^{2}u_{2}^{2}-m_{1}^{2}m_{2}^{2})-2e^{-t/2}u_{1}u_{2}{\rm Re}[v(z_{1})\overline{z_{2}}]\Big{]}-\kappa_{4}\partial_{\eta_{1}}\partial_{\eta_{2}}m_{1}^{2}m_{2}^{2}+\operatorname{O}(N^{-c}),\end{split}$ (5.9)
where we used the notations $m_{1}:=m^{v(z_{1})}(\mathrm{i}\eta_{1})$, $u_{1}:=u^{v(z_{1})}(\mathrm{i}\eta_{1})$, and similar notations for $m_{2},u_{2}$.
@@
$\begin{split}\mathbb{E}I_{\eta_{c}}^{T}(f,t)I_{\eta_{c}}^{T}(g,0)&=\frac{1}{8\pi^{2}}\int\int\Delta f(z_{1})\Delta g(z_{2})\,{\mathrm{d}}z_{1}{\mathrm{d}}z_{2}\\ &\quad\times\bigg{[}-\frac{1}{2}\log\bigg{[}1+|z_{2}v(z_{1})|^{2}(u^{v(z_{1})}(\mathrm{i}\eta_{0}(\eta_{1})))^{2}(u^{z_{2}}(\mathrm{i}\eta_{2}))^{2}\\ &\qquad\quad-(m^{v(z_{1})}(\mathrm{i}\eta_{0}(\eta_{1})))^{2}(m^{z_{2}}(\mathrm{i}\eta_{2}))^{2}-2u^{v(z_{1})}(\mathrm{i}\eta_{0}(\eta_{1}))u^{z_{2}}(\mathrm{i}\eta_{2}){\rm Re}[v(z_{1})\overline{z_{2}}]\bigg{]}\\ &\qquad\quad+\frac{\kappa_{4}}{2}m^{v(z_{1})}(\mathrm{i}\eta_{0}(\eta_{1}))^{2}m^{z_{2}}(\mathrm{i}\eta_{2})^{2}\bigg{]}\bigg{|}_{\eta_{1}=0,\atop\eta_{2}=0}+\operatorname{O}(N^{-c}).\end{split}$ (5.10)
@@
$\begin{split}&\log\big{[}1+|z_{2}v(z_{1})|^{2}(u^{v(z_{1})}(\mathrm{i}\eta_{0}(\eta_{1})))^{2}(u^{z_{2}}(\mathrm{i}\eta_{2}))^{2}-(m^{v(z_{1})}(\mathrm{i}\eta_{0}(\eta_{1})))^{2}(m^{z_{2}}(\mathrm{i}\eta_{2}))^{2}-2u^{v(z_{1})}(\mathrm{i}\eta_{0}(\eta_{1}))u^{z_{2}}(\mathrm{i}\eta_{2}){\rm Re}[v(z_{1})\overline{z_{2}}]\big{]}\\ &\quad=\log\big{[}1+e^{-t}(|z_{2}z_{1}|^{2}u^{z_{1}}(i\eta_{1})^{2}(u^{z_{2}}(\mathrm{i}\eta_{2}))^{2}-m^{z_{1}}(i\eta_{1})^{2}(m^{z_{2}}(\mathrm{i}\eta_{2}))^{2})-2e^{-t/2}u^{z_{1}}(i\eta_{1})u^{z_{2}}(\mathrm{i}\eta_{2}){\rm Re}[z_{1}\overline{z_{2}}]\big{]}.\end{split}$ (5.11)
@@
$m^{v(z_{1})}(\mathrm{i}\eta_{0}(\eta_{1}))^{2}=e^{-t}m^{z_{1}}(\mathrm{i}\eta_{1})=e^{-t}(1-|z_{1}|^{2}),\qquad m^{z_{2}}(\mathrm{i}\eta_{2})^{2}=(1-|z_{2}|^{2}).$ (5.13)
@@
We now compute the leading deterministic term in (5.18). For this purpose we recall the multi–resolvent local law from [CipErdSch2022bis, Theorem 3.3]. Define $G_{i}:=(W-Z_{i}-\mathrm{i}\eta_{i})^{-1}$, with $Z_{i}$ as in (3.3), and denote by $M_{i}$ its deterministic approximation. Then, for a deterministic matrix $A\in\mathbb{C}^{2N\times 2N}$, the deterministic approximation of $G_{1}AG_{2}$ is given by
@@
By the local law from [CipErdSch2022bis, Theorem 3.3], we then have
@@
$\begin{split}&\mathbb{E}|I_{2}|^{2}=\mathbb{E}\left|\frac{1}{4\pi\mathrm{i}\sqrt{N}}\int_{\mathbb{C}}\Delta f(z)\int_{0}^{t}\int_{\eta_{s}(\eta_{c},t)}^{\eta_{s}(T,t)}\mathrm{Tr}\big{[}G_{s}^{z_{s}(z)}(\mathrm{i}\eta^{\prime})^{2}E_{i}{\mathrm{d}}B_{s}E_{j}\big{]}\Bigg{]}\,{\mathrm{d}}\eta^{\prime}{\mathrm{d}}z^{2}\right|^{2}\\ &\quad=-\frac{1}{16\pi^{2}}\int_{\mathbb{C}}\int_{\mathbb{C}}\Delta f(z_{1})\Delta f(z_{2})\times\int_{0}^{t}\partial s\begin{cases}\log\big{[}|(1-e^{-2(t-s)})(1-|z_{1}|^{2})+|z_{1}-e^{-(t-s)}z_{2}|^{2}\big{]}&|z_{1}|,|z_{2}|\leqslant 1\\ \log|z_{l}-e^{-(t-s)}z_{m}|^{2}-\log|z_{l}|^{2}&|z_{m}|\leqslant 1,\,|z_{l}|>1\\ \log|e^{-(t-s)}-z_{1}\overline{z_{2}}|^{2}-\log|z_{1}z_{2}|^{2}&|z_{1}|,|z_{2}|>1\end{cases}\\ &\quad=-\frac{1}{16\pi^{2}}\int_{\mathbb{C}}\int_{\mathbb{C}}\Delta f(z_{1})\Delta f(z_{2})\times\begin{cases}\log|z_{1}-z_{2}|^{2}-\log\big{[}|(1-e^{-2t)})(1-|z_{1}|^{2})+|z_{1}-e^{-t}z_{2}|^{2}\big{]}&|z_{1}|,|z_{2}|\leqslant 1\\ \log|z_{1}-z_{2}|^{2}-\log|z_{l}-e^{-t}z_{m}|^{2}&|z_{m}|\leqslant 1,\,|z_{l}|>1\\ \log|1-z_{1}\overline{z_{2}}|^{2}-\log|e^{-t}-z_{1}\overline{z_{2}}|^{2}&|z_{1}|,|z_{2}|>1\\ \end{cases}\end{split}$ (5.21)

### v2

@@ 4 Independence of singular values
${\rm d}f_{s}=m_{s}(z)\partial_{z}f_{s}{\rm d}s+\frac{1}{2N}\sum_{1\leqslant|k|\leqslant N}\frac{v_{k}(s)({\rm d}\langle b_{k}\rangle_{s}-{\rm d}s)}{(x_{k}(s)-z)^{3}}-\frac{1}{\sqrt{2N}}\sum_{1\leqslant|k|\leqslant N}\frac{v_{k}(s)}{(z-x_{k}(s))^{2}}{\rm d}b_{k}(s).$ (4.72)
@@
$\begin{split}{\rm d}g_{u\wedge\tau}(z)=(m_{u}(z_{t-u})-\widetilde{m}_{u}(z_{t-u}))\partial_{z}f_{u}(z_{t-u}){\rm d}({u\wedge\tau})+\frac{1}{2N}\sum_{1\leqslant|i|\leqslant N}\frac{({\rm d}\langle b_{i}\rangle_{u}/{\rm d}u-1)v_{i}(u)}{(x_{i}(u)-z_{t-u})^{3}}{\rm d}(u\wedge\tau)\\ -\frac{\operatorname{O}(1)}{\sqrt{N}}\sum_{1\leqslant|i|\leqslant N}\frac{{v}_{i}({u})}{(z_{t-u}-x_{i}(u))^{2}}{\rm d}b_{i}({u\wedge\tau}).\end{split}$ (4.74)
@@
where the above integral is due to points $x_{i}(u)\geqslant R/10$, and use the local law in the domain $R/10<x_{i}<R/2$, the upper bound $x_{i}<C$, and the maximum principle bound $|v_{i}|\leqslant 1$. The above integral contributes to at most
@@
$\frac{1}{N}\sum_{1\leqslant|i|\leqslant N}\frac{({\rm d}\langle b_{i}\rangle_{u}/{\rm d}u-1)v_{i}(u)}{(x_{i}(u)-z_{t-u})^{3}}=\partial_{zz}\frac{1}{2N}\sum_{1\leqslant|i|\leqslant N}\frac{({\rm d}\langle b_{i}\rangle_{u}/{\rm d}u-1)v_{i}(u)}{x_{i}(u)-z_{t-u}}=\frac{1}{N}\sum_{1\leqslant|i|\leqslant N}\partial_{zz}\frac{\operatorname{O}(1)x_{i}v_{i}(u)}{x_{i}(u)^{2}-z_{t-u}^{2}}.$ (4.79)
@@
$M_{s}$ $:=\operatorname{Im}\int_{0}^{s}\frac{1}{\sqrt{N}}\sum_{1\leqslant|k|\leqslant N}\frac{{v}_{k}({u})}{(z_{t-u}-x_{k}(u))^{2}}{\rm d}b_{k}(u\wedge\tau)$
@@
$\sup_{0\leqslant s\leqslant t}|M_{s}|^{2}\leqslant\varphi^{1/10}\int_{0}^{t}\frac{\eta_{t-u}^{2}}{N}\left(\sum_{k=1}^{N}\left|\frac{E-x_{k}(u)}{|z_{t-u}-x_{k}(u)|^{4}}+\frac{E+x_{k}(u)}{|z_{t-u}+x_{k}(u)|^{4}}\right|{v}_{k}({u})\right)^{2}{\rm d}(u\wedge\tau).$ (4.80)
With (4.78) and the local law ($\tau\leqslant\tau_{0}$) the above sum is bounded with (we omit the terms corresponding to $x_{i}>R/10$ as they are shown to be negligible thanks to the local law and the simple bound $|v_{i}|\leqslant 1$)
$\phantom{{}={}}N^{1/2}\int_{0}^{c}\left|\frac{E-x}{|z_{t-u}-x|^{4}}+\frac{E+x}{|z_{t-u}+x|^{4}}\right|\frac{x}{\sqrt{\max(x,u)}}{\rm d}x+\varphi N\int_{0}^{c}\left|\frac{E-x}{|z_{t-u}-x|^{4}}+\frac{E+x}{|z_{t-u}+x|^{4}}\right|x\max(x,u){\rm d}x$
@@
For any $t\in[\varphi^{100}/N,1]$ and $1\leqslant i<N$, uniformly in $\alpha$ we have, with overwhelming probability,
@@
We have $|v_{i}(0)|\leqslant\varphi(v_{i}^{(1)}(0)+v_{i}^{(2)}(0))$ from (4.64); from Lemma 4.20 (i), this implies that for any $t\geqslant 0,1\leqslant i\leqslant N$ we have $|v_{i}(t)|\leqslant\varphi(v_{i}^{(1)}(t)+v_{i}^{(2)}(t))$. From Equation (4.77) and Lemmas 4.22 and 4.23, we have
@@ 5 Resolvent fluctuations from the dynamics noise
$\frac{1}{m^{z}}=\mathrm{i}\eta+m^{z}-\frac{|z|^{2}}{\mathrm{i}\eta+m^{z}},\qquad\eta{\rm Im}[m^{z}]>0.$
@@
In this section we analyze the evolution of the resolvent $G_{t}^{z}(\mathrm{i}\eta):=(W_{t}-Z-\mathrm{i}\eta)^{-1}$ along the flow (1.5). In particular, we will study the evolution of the resolvent along the characteristics [^7] Our convention here is that the characteristics move upwards as in (4.40), contrary to (4.69). (cf. [CipErdSch2022bis, Eq. (5.4)] and [CipErdXu2023bis, Eq. (5.4)]):
$\partial_{s}\eta_{s}=-{\rm Im}m^{z_{s}}(\mathrm{i}\eta_{s})-\frac{\eta_{s}}{2},\qquad\quad\partial_{s}z_{s}=-\frac{z_{s}}{2}.$ (5.1)
In the following, for $\eta_{*}>0$ and $z_{*}\in\mathbb{C}$, by $\eta_{s}(\eta_{*},t),z_{s}(z_{*})$, for $s\in[0,t]$, we denote the solutions of (5.1) such that at time $t$ satisfy $\eta_{t}(\eta_{*},t)=\eta_{*}$ and $z_{t}(z_{*})=z_{*}$.
Here we perform the analysis in the regime $|z|\leqslant 1-N^{-\varepsilon}$, for an arbitrary small $\varepsilon>0$, since this is the regime of interest of Proposition 3.3; the regime $|z|>1-N^{-\varepsilon}$ is negligible (see e.g. above [CipErdSch2023, Eq. (4.25)]). The main result of this section is a decomposition theorem for $I_{\eta_{c}}^{T}(f,t)$ into the sum of two terms: one that depends only on the initial condition and one that is a martingale (see Proposition 5.1). Before stating this result we introduce the notation
@@
In order to control the evolution of the resolvent along the flow (5.4) we consider the characteristics (5.1). Note that along the characteristics we have
@@
with very high probability, where we used that $|\langle G_{t}^{2}-M_{t}^{\prime}\rangle|\lesssim N^{\xi}(N\eta_{t}^{2})^{-1}$ for any arbitrary small $\xi>0$ by [CipErdXu2025, Eq. (3.26)] for $z_{1}=z_{2}$ and $\eta_{1}=\eta_{2}$, and to estimate the error term we used that for $s_{1}\leqslant s_{2}\leqslant t$ we have
$\exp\left(\int_{s_{1}}^{s_{2}}\left[\frac{1}{2}+\partial_{\eta_{\tau}}m^{z_{\tau}}(\mathrm{i}\eta_{\tau})\right]\,{\mathrm{d}}\tau\right)\lesssim\frac{\eta_{s_{1}}}{\eta_{s_{2}}},$
as a consequence of $|\partial_{\eta_{\tau}}m^{z_{\tau}}(\mathrm{i}\eta_{\tau})|\leqslant\rho_{\tau}/\eta_{\tau}$.
@@
Before presenting these proofs we comment on their relation with the proof of [CipErdSch2023, Proposition 3.3]. In the reminder of this section the local law [CipErdSch2022bis, Theorem 5.2] is a fundamental input to compute the deterministic approximation of product of resolvents of the form $G_{s}^{z_{1}}G_{s}^{z_{2}}$, i.e. this local law is used for any fixed time $0\leqslant s\leqslant t$, but only for resolvents evaluated at the same time. On the other hand, the CLT for resolvents [CipErdSch2023, Proposition 3.3] is used only at time $s=0$ (see (5.9) below) to compute the correlation of the initial condition.
@@
From now on $c>0$ is a small fixed constant that may change from line to line. By [CipErdSch2023, Proposition 3.3], for $\eta_{i}\geqslant n^{-1+\delta_{1}}$, it follows that [^8] The term $|z_{1}-z_{2}|^{-4}$ in [CipErdSch2023, Eq. (3.13)] is just a consequence of the sub-optimal local law in [CipErdSch2023, Eq. (5.9)], and it can readily be removed using the local law [CipErdXu2025, Eq. (3.26)] (see e.g. [CipErdSch2022bis, Proposition 3.4] for the regime $|z|\leqslant 1-c$). Additionally, if $z_{0}(z_{1})>1-N^{-C\varepsilon}$, for a fixed $C>0$, then inspecting the proof of [CipErdSch2023, Proposition 3.3] it is easy to see that the term $1-|z_{i}|$ in [CipErdSch2023, Eq. (3.13)] can be replaced by $N^{C\varepsilon}$. In the complementary regime we use [CipErdSch2023, Proposition 3.3] verbatim.
$\begin{split}&\mathbb{E}\mathrm{Tr}\big[G_{0}^{z_{0}(z_{1})}(\mathrm{i}\eta_{1})-\mathbb{E}G_{0}^{z_{0}(z_{1})}(\mathrm{i}\eta_{1})\big]\mathrm{Tr}\big[G_{0}^{z_{2}}(\mathrm{i}\eta_{2})-\mathbb{E}G_{0}^{z_{2}}(\mathrm{i}\eta_{2})\big]\\ &=\partial_{\eta_{1}}\partial_{\eta_{2}}\log\Big[1+e^{-t}(|z_{0}(z_{1})z_{2}|^{2}u_{1}^{2}u_{2}^{2}-m_{1}^{2}m_{2}^{2})-2e^{-t/2}u_{1}u_{2}{\rm Re}[z_{0}(z_{1})\overline{z_{2}}]\Big]-\kappa_{4}\partial_{\eta_{1}}\partial_{\eta_{2}}m_{1}^{2}m_{2}^{2}+\operatorname{O}\left(\frac{N^{-c}}{\eta_{1}\eta_{2}}\right),\end{split}$ (5.9)
where we used the notations $m_{1}:=m^{z_{0}(z_{1})}(\mathrm{i}\eta_{1})$, $u_{1}:=u^{z_{0}(z_{1})}(\mathrm{i}\eta_{1})$, and similar notations for $m_{2},u_{2}$.
@@
$\begin{split}\mathbb{E}I_{\eta_{c}}^{T}(f,t)I_{\eta_{c}}^{T}(g,0)&=\frac{1}{8\pi^{2}}\int\int\Delta f(z_{1})\Delta g(z_{2})\,{\mathrm{d}}z_{1}{\mathrm{d}}z_{2}\\ &\quad\times\bigg[-\frac{1}{2}\log\bigg[1+|z_{2}z_{0}(z_{1})|^{2}(u^{z_{0}(z_{1})}(\mathrm{i}\eta_{0}(\eta_{1})))^{2}(u^{z_{2}}(\mathrm{i}\eta_{2}))^{2}\\ &\qquad\quad-(m^{z_{0}(z_{1})}(\mathrm{i}\eta_{0}(\eta_{1})))^{2}(m^{z_{2}}(\mathrm{i}\eta_{2}))^{2}-2u^{z_{0}(z_{1})}(\mathrm{i}\eta_{0}(\eta_{1}))u^{z_{2}}(\mathrm{i}\eta_{2}){\rm Re}[z_{0}(z_{1})\overline{z_{2}}]\bigg]\\ &\qquad\quad+\frac{\kappa_{4}}{2}m^{z_{0}(z_{1})}(\mathrm{i}\eta_{0}(\eta_{1}))^{2}m^{z_{2}}(\mathrm{i}\eta_{2})^{2}\bigg]\bigg|_{\eta_{1}=0,\atop\eta_{2}=0}+\operatorname{O}(N^{-c}).\end{split}$ (5.10)
@@
$\begin{split}&\log\big[1+|z_{2}z_{0}(z_{1})|^{2}(u^{z_{0}(z_{1})}(\mathrm{i}\eta_{0}(\eta_{1})))^{2}(u^{z_{2}}(\mathrm{i}\eta_{2}))^{2}-(m^{z_{0}(z_{1})}(\mathrm{i}\eta_{0}(\eta_{1})))^{2}(m^{z_{2}}(\mathrm{i}\eta_{2}))^{2}-2u^{z_{0}(z_{1})}(\mathrm{i}\eta_{0}(\eta_{1}))u^{z_{2}}(\mathrm{i}\eta_{2}){\rm Re}[z_{0}(z_{1})\overline{z_{2}}]\big]\\ &\quad=\log\big[1+e^{-t}(|z_{2}z_{1}|^{2}u^{z_{1}}(i\eta_{1})^{2}(u^{z_{2}}(\mathrm{i}\eta_{2}))^{2}-m^{z_{1}}(i\eta_{1})^{2}(m^{z_{2}}(\mathrm{i}\eta_{2}))^{2})-2e^{-t/2}u^{z_{1}}(i\eta_{1})u^{z_{2}}(\mathrm{i}\eta_{2}){\rm Re}[z_{1}\overline{z_{2}}]\big].\end{split}$ (5.11)
@@
$m^{z_{0}(z_{1})}(\mathrm{i}\eta_{0}(\eta_{1}))^{2}=e^{-t}m^{z_{1}}(\mathrm{i}\eta_{1})=e^{-t}(1-|z_{1}|^{2}),\qquad m^{z_{2}}(\mathrm{i}\eta_{2})^{2}=(1-|z_{2}|^{2}).$ (5.13)
@@
We now compute the leading deterministic term in (5.18). For this purpose we recall the multi–resolvent local law from [CipErdXu2025, Eq. (3.26)]. Define $G_{i}:=(W-Z_{i}-\mathrm{i}\eta_{i})^{-1}$, with $Z_{i}$ as in (3.3), and denote by $M_{i}$ its deterministic approximation. Then, for a deterministic matrix $A\in\mathbb{C}^{2N\times 2N}$, the deterministic approximation of $G_{1}AG_{2}$ is given by
@@
By the local law from [CipErdXu2025, Eq. (3.26)], we then have
@@
[…]